Copper sulphate solution is electolysed using platinum electrodes A. current of 0.193 amperes is passed for 3hrs. How many grams of copper are deposited? (cu = 63.5, F = 96500 coulombs)
1
Expert's answer
2021-05-09T00:01:03-0400

Time = "2hrs. = 120 \times 60 = 7200s"


Current "= 0.193A"


Quantity of electricity "= 120 \times 0.193 = 60 \times 23.16"


96,500 deposits 63.5 gm of Cu.


60 x 23.2 will deposit "\dfrac{63.5 \times 60 \times 23.2 }{96500} = 0.457g" of Copper.